OpenStudy (anonymous):

@ phi @ Wio @mathmale @IMStuck @ShailKumar Tyler has a rectangular garden that measures 10 m wide by 13 m long. He wants to increase the area to 208 m2 by increasing the width and length by the same amount. What will be the length (longer dimension) of the new garden? A. 14 m wide B. 15 m wide C. 16 m wide D. 17 m wide

OpenStudy (anonymous):

i think c

OpenStudy (anonymous):

@Zale101

OpenStudy (imstuck):

In this one you have to solve for x in order to find the new dimensions. 208 = (10 + x)(13 + x). You have to multiply this out (FOIL it) in order to find a p[olynomial to factor.

OpenStudy (anonymous):

please show me if you can IMstuck

OpenStudy (anonymous):

or sarah

OpenStudy (anonymous):

The answer is C. Since Tyler wants to increase the area to 208 meters, he will have to increase the width and length. Whatever he does to the width, he has to do to the length. Therefore, if he adds 3 meters to the width, he will have to add 3 meters to the length. (10+x)(13+x)=208 (10+3)(13+3)=208 (13)(16)=208 The answer is C because if you add 3 meters to the length and 3 meters to the width, the new measures will be 13 and 16. Therefore, if you multiply these two together, you will get 208 meters.

OpenStudy (anonymous):

thats what i was thinking

OpenStudy (anonymous):

ill wait for IMstuck to replie to pive medal

OpenStudy (anonymous):

just for fareness

OpenStudy (imstuck):

Ok when you foil out you get x^2 + 23x + 130 = 208. You have to move the 208 over to the side with all the other numbers and set the polynomial equal to zero.\[x ^{2}+23x+130-208=0\]Doing the math you get \[x ^{2}+23x-78=0\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im fallowing

OpenStudy (anonymous):

sorry IMstuck but

OpenStudy (anonymous):

sarah plz give IMstuck a medal

OpenStudy (anonymous):

for fairness

OpenStudy (imstuck):

Now you have to factor this out. I'm trying to find the factors of 78 right now. It's a tricky number so I'm going to do the quadratic formula on it. BRB. Just take a sec.

OpenStudy (anonymous):

ok im going to submit it wish me luck

OpenStudy (anonymous):

any objections?

OpenStudy (imstuck):

Ok, the factors are (x - 3)(x + 3). But of course as you know, a length of something cannot be negative, so the answer is x = 3. So the length of the longer dimension is 13 + x or 13 + 3 = 16.

OpenStudy (anonymous):

did it

OpenStudy (imstuck):

Did you get 3?

OpenStudy (anonymous):

100%!!! thank you!!

OpenStudy (anonymous):

yes i did

OpenStudy (imstuck):

Yay for 100%'s!

OpenStudy (anonymous):

you sir have a fan

OpenStudy (anonymous):

IMstuck

OpenStudy (imstuck):

Actually, I am a "ma'am" not a "sir", but that's ok! Thanks!

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

i did not know

OpenStudy (imstuck):

I wouldn't expect you to! I do not know if you are male or female by your name here either!

OpenStudy (anonymous):

im a guy

OpenStudy (imstuck):

Good luck with your math stuff! I'm a high school math teacher and love to see students so concerned with their success in math.

OpenStudy (anonymous):

btw you have to say my name as tow thousand (pause for half a sec) twenty eight

OpenStudy (anonymous):

and thank you thank you again

OpenStudy (anonymous):

*two

OpenStudy (imstuck):

You are so very very welcome!!!

OpenStudy (anonymous):

i might do one more math tonight but im not sure

OpenStudy (imstuck):

One more math what? One more problem?

OpenStudy (anonymous):

lesson

OpenStudy (anonymous):

ill tag you if i do ok bye for now

OpenStudy (imstuck):

Gotcha. If you need help... : )